OpenStudy (anonymous):

solve for x : −5|x + 1| = 10

OpenStudy (ageta):

Let's solve your equation step-by-step. −5(|x+1|)=10 Step 1: Divide both sides by -5. −5(|x+1|)−5=10−5 |x+1|=−2 Step 2: Solve Absolute Value. |x+1|=−2 No solutions. (Absolute value cannot be less than 0.)

OpenStudy (ageta):

No solutions.

OpenStudy (anonymous):

ok, thank you so much , that really helped

OpenStudy (ageta):

your welcome

OpenStudy (tkhunny):

Well, I guess you can waste your time on this one. For me, this is an "Eyeball Problem". Just look at it and write down the answer. -5|x + 1| = 10 10 > 0 - Point #1 |x+1| > 0 - That's what absolute values do. -5|x+1| < 0 - Point #2 How many numbers can you name which are both greater than zero (point #1) and less than zero (point #2)? If you just reach out and grab problems like this, you can have more time to work on harder problems.